find an equation of the line passing through the given points, and write the equation in both standard and slope-intercept form (if possible)
(3/4, 8/3) and (2/5, 2/3)

Answer by mananth(16946)

x1 y1 x2 y2
3/ 4 8/3 2/5 2/3

slope m = (y2-y1)/(x2-x1)
( 2/3 - 8/3 )/( 2/5 - 3/4 )
( -2 / -7/20 )
m= 40/7

Plug value of the slope and point ( 3/4, 8/3 ) in
Y = m x + b
8/3 = 30/7 + b
b= 8/3 +30/7
b= 26/21
So the equation will be
Y = (40/7)x +26/21
multiply by 21
21y = 120x+26
-120x+21y==26

m.ananth@hotmail.ca